OpenStudy (anonymous):

A box has dimensions of 19 inches long, 1.7 feet wide, and 6 inches high. What is the volume of the box? The formula for the volume is V = l • w • h. (1 point) 193.8 cubic inches (in3) 2325.6 cubic inches (in3) 1.35 cubic inches (in3) 16.15 cubic inches (in3)

OpenStudy (anonymous):

let l, b and h be the length width and height of the box therefore l=19 inches b=1.7 feet =1.7*12=20.4 inches h= 6 inches Now volume of the box will be given as V= l*b*h= 19*20.4*6=2325.6 cubic inches (in^3)

OpenStudy (anonymous):

@juleenelson

OpenStudy (anonymous):

thankyou!:) im no math genius...lol:p :)

OpenStudy (anonymous):

@juleenelson Welcome